# Cross-Modal Consistency Learning for Sign Language Recognition

### v1

@@ 3.3 Cross-Modal Contrastive Learning
While previous work [49] successfully employed Kullback-Leibler (KL) divergence to maintain distribution consistency within pose modality, cross-modal alignment between pose and RGB modality remains challenging due to their substantial data distribution gap. To this end, we extend single-modal contrastive learning to the cross-modal domain by directly aligning embeddings from both modalities in the latent space. Specifically, for RGB features $f^{R}_{q}$, we utilize $f^{P}_{k}$ as positive samples and embeddings in $\mathcal{M}^{P}_{f}$ as negative samples from the pose modality to perform contrastive learning; analogously, the symmetric procedure is applied to pose features. The bidirectional cross-modal contrastive losses are defined as:
@@ 3.4 Model Details
During pre-training phase, each projection head is a 2-layer MLP which projcets the encoder output into a 128-dimensional embedding. During fine-tuning phase, we utilize the query encoders as feature extractors and replace the projection head with a fully connected layer for prediction. In this phase, both branches are optimized via cross-entropy loss. The final prediction is obtained by summing the predicted logits from the RGB and pose branches.
@@ 4.1 Datasets and Evaluation Metrics
MSASL is an American sign language (ASL) dataset, comprising 25,513 samples with a vocabulary size of 1,000. It consists of 16,054, 5,287, and 4,172 samples in the training, development (dev), and test set, respectively. Furthermore, it provides subset consisting of only the top 200 most frequent glosses. WLASL is another widely-used ASL dataset with a vocabulary size of 2000. It includes 14,289, 3,916, and 2,878 samples in the training, dev, and test set, respectively. Similar to MSASL, WLASL also releases subset consisting of 300 frequent glosses.
NMFs-CSL is a challenging Chinese sign language (CSL) dataset that incorporates a wide range of fine-grained non-manual features (NMFs). With a vocabulary size of 1,067, the dataset has 25,608 and 6,402 samples in the training and test sets. SLR500 is another large-scale CSL dataset that covers 500 common words. It contains a total of 125,000 samples, with 90,000 samples for training and 35,000 samples for testing. Following [13, 55], we report the average accuracy over instances and classes, respectively. Since NMFs-CSL and SLR500 have a balanced distribution with an equal number of samples per class, we only report per-instance accuracy.
4.2 Experimental Settings
The RGB branch applies spatial augmentations including random cropping, horizontal flipping, and color jittering, while the pose branch implements random rotation, scaling, joint masking, and flipping. Both branches share temporal augmentation through random temporal cropping. More details about data preprocessing, pose encoder and data augmentation are provided in the Appendix.
The temperature hyperparameters $\tau$ and $\tau_{spm}$ are set to $0.07$ and $0.02$ respectively. The hyperparameters of MPM $\alpha$ and $p$ are set to $0.5$, and $0.2$, respectively. During pre-training phase, CCL-SLR is trained for 140 epochs with a learning rate of 0.01 and batch size 64, where the learning rate is divided by 10 at epoch 100. During fine-tuning phase, the model is trained for 40 epochs with learning rate 0.05 and batch size 128, with the learning rate divided by 10 at epochs 25 and 35. An SGD optimizer with 0.9 momentum is utilized throughout the training process. All experiments are conducted on 8$\times$ NVIDIA RTX 3090 GPUs.
@@ 4.3 Comparison with State-of-the-art Methods
NMFs-CSL. The comparison on the NMFs-CSL dataset is shown in Table 2. Unlike previous pre-training methods such as SignBERT[13] and BEST[46], which simply fuse RGB predictions in downstream tasks and lack consideration of modeling cross-modal consistency during pre-training. Our method achieves superior performance with 6.0% improvements over SignBERT on NMFs-CSL dataset.
SLR500. Lastly, Table 2 also presents the performance comparison on the SLR500 dataset. On this sufficiently large dataset, self-supervised learning methods such as SignBERT+[16] and BEST[46] achieve more outstanding performance. Similarly, our cross-modal pre-training approach effectively integrates consistency cues from two modalities, achieving a 97.8% Top-1 accuracy.
@@
Table 3: Ablation studies for the major components of CCL-SLR.
@@ 4.4 Ablation Study
In this section, we conduct several ablation studies to verify the effectiveness of several key components in our proposed framework on the MSASL [48] dataset.
@@
Cross-Modal Contrastive Learning Loss. As shown in Table 4, we first study the effects of the cross-modal contrastive learning loss $\mathcal{L}_{cross}$. After removing $\mathcal{L}_{cross}$ from CCL-SLR, the top-1 accuracy significantly drops by about 5%. The reason is that $\mathcal{L}_{cross}$ plays a significantly important role in our framework by explicitly and effectively aligning the feature spaces between the two modalities, which implicitly helps the model extract cross-modal consistency cues. Additionally, Fig. 4 demonstrates the heatmaps generated by fine-tuning RGB encoder from pre-training with and without $\mathcal{L}_{cross}$. Especially, performing cross-modal contrastive learning during pretraining allows the RGB branch to learn semantic information from the pose modality. This significantly enhances the model’s attention to non-manual features (yellow and red areas) such as facial expressions and body movements.
Pre-Training Data Scale. In Fig. 5, we explore the impact of the pre-training data scale on the performance of our proposed framework. The experimental results reveal that the performance exhibits a steady improvement as the proportion of the pre-training data scale increases, demonstrating that our model can effectively leverage large-scale unlabeled multi-modal sign language data to enhance its representation capability, thereby indicating the scalability of our approach to large-scale sign language data.
Figure 4: Visualizations of heatmaps by Grad-CAM [35]. Top: raw frames; Middle: heatmaps of the fine-tuning results of the RGB encoder from pre-training without $\mathcal{L}_{cross}$; Bottom: heatmaps of the fine-tuning results of the RGB encoder from pre-training with $\mathcal{L}_{cross}$. The samples are from the NMFs-CSL [15] dataset.
Figure 5: Hyperparameter selection on CCL-SLR performance. We sequentially investigate the impact of the pre-training data scale, the number of pseudo-positive labels $k$, and the mask probability $\alpha$.
Motion-Preserving Masking. The influence of MPM is explored by sequentially changing the mask probability parameter $\alpha$ within $[0.0,1.0]$. As shown in Fig.5, the performance reaches its optimum at $\alpha=0.2$, suggesting that moderate masking effectively mitigate static bias while preserving essential dynamic information. However, an excessively high mask ratio can lead to performance degradation. Visualization results of MPM are provided in the Appendix.
Semantic Positive Mining. We investigate the sensitivity of hyperparameter $k$ to our method, which denotes the number of mined semantic positives per modality. As illustrated in Fig. 5, experimental results demonstrate that either extremely small or large values of $k$ lead to performance degradation. Excessively large k value will introduce noise during contrastive learning, which affects the representation capability of the model. According to these empirical results, we determine $k=10$ as default.
@@
Table 6: Ablation studies for different pre-training modalities.
Figure 6: t-SNE visualizations of feature embeddings. We sample 30 sign words from SLR500 dataset and visualize the features extracted from the proposed CCL-SLR and its baseline, respectively.
Why keeping two memory banks? Different from vanilla MoCo v2 [3] framework, we maintain two memory banks ($\mathcal{M}_{g}$ and $\mathcal{M}_{f}$) in our pipeline: one for pseudo-labeling and the other for contrastive learning. Intuitively, maintaining a unique memory bank is structurally more concise. However, according to previous work [9], features before projection are more generally applicable, while features after projection are specialized for the instance discrimination task of contrastive learning. As shown in the Table 5, experimental results also indicate that directly using $\mathcal{M}_{f}$ for pseudo-labeling leads to a slight performance degradation.
@@ 5 Conclusion
In this paper, we investigate the challenge of sign language recognition through multi-modal consistency learning. We propose CCL-SLR, a novel self-supervised pre-training framework that enhances sign representation by leveraging complementary information across modalities. Our framework incorporates two key innovations: MPM, which enables motion-aware feature extraction from RGB videos, and SPM, which effectively addresses the semantic collision problem in contrastive learning. We conduct comprehensive evaluations on four widely-used ISLR datasets, where CCL-SLR consistently achieves state-of-the-art performance, validating the effectiveness of our proposed approach.

### v2

@@ 3.3 Cross-Modal Contrastive Learning
While previous work [49] successfully employed Kullback-Leibler (KL) divergence to ensure distribution consistency within the pose modality, cross-modal alignment between pose and RGB modality remains challenging due to the substantial distribution gap between them. To this end, we extend single-modal contrastive learning to the cross-modal domain by directly aligning embeddings from both modalities in the latent space. Specifically, for RGB features $f^{R}_{q}$, we utilize $f^{P}_{k}$ as positive samples while treating embeddings in $\mathcal{M}^{P}_{f}$ as negative. A symmetric procedure is applied to pose features. The bidirectional cross-modal contrastive loss functions are defined as:
@@ 3.4 Model Details
During the pre-training phase, each projection head is a 2-layer MLP which projects the encoder outputs into a 128-dimensional embedding. During the fine-tuning phase, we employ the query encoders as feature extractors and replace the projection head with a fully connected layer for prediction. In this phase, both branches are optimized via cross-entropy loss. The final prediction is obtained by summing the predicted logits from the RGB and pose branches.
@@ 4.1 Datasets and Evaluation Metrics
MSASL is an American Sign Language (ASL) dataset, consisting of 25,513 samples with a vocabulary size of 1,000. It includes 16,054, 5,287, and 4,172 samples in the training, development (dev), and test set, respectively. Additionally, it provides a subset containing only the top 200 most frequent glosses. WLASL is another widely used ASL dataset with a vocabulary size of 2,000. It includes 14,289, 3,916, and 2,878 samples in the training, dev, and test sets, respectively. Similar to MSASL, WLASL also provides a subset of 300 frequent glosses.
NMFs-CSL is a challenging Chinese Sign Language (CSL) dataset that incorporates a wide range of fine-grained non-manual features (NMFs). With a vocabulary size of 1,067, it contains 25,608 training samples and 6,402 test samples. SLR500 is another large-scale CSL dataset covering 500 common words. It contains a total of 125,000 samples, with 90,000 samples for training and 35,000 samples for testing. Following [13, 55], we report the average accuracy over instances and classes, respectively. Since NMFs-CSL and SLR500 exhibit a balanced distribution with an equal number of samples per class, we only report per-instance accuracy.
4.2 Implementation Details
The RGB branch applies spatial augmentations including random cropping, horizontal flipping, and color jittering, while the pose branch implements random rotation, scaling, joint masking, and flipping. Both branches share temporal augmentation through random temporal cropping. More details about data preprocessing, the pose encoder and data augmentation are provided in the Appendix.
The temperature hyperparameters $\tau$ and $\tau_{spm}$ are set to $0.07$ and $0.02$, respectively. The hyperparameters of MPM $p$ and $\alpha$ are set to $0.5$ and $0.2$, respectively. During the pre-training phase, CCL-SLR is trained for 140 epochs with a learning rate of 0.01 and batch size 64, with the learning rate reduced by a factor of 10 at epoch 100. During the fine-tuning phase, the model is trained for 40 epochs with a learning rate of 0.05 and a batch size of 128, with the learning rate reduced by a factor of 10 at epochs 25 and 35. A Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 is utilized throughout the training process. All experiments are conducted on 8$\times$ NVIDIA RTX 3090 GPUs.
@@ 4.3 Comparison with State-of-the-art Methods
NMFs-CSL. The comparison on the NMFs-CSL dataset is shown in Table 2. Unlike previous pre-training methods such as SignBERT[13] or BEST[46], which simply fuse RGB predictions in downstream tasks and lack consideration of modeling cross-modal consistency during pre-training. Our method achieves superior performance with 6.0% and 5.2% improvements over SignBERT and BEST on the NMFs-CSL dataset, respectively.
SLR500. Lastly, Table 2 also presents the performance comparison on the SLR500 dataset. On this sufficiently large dataset, self-supervised learning methods such as SignBERT+[16] and BEST[46] achieve more outstanding performance. Similarly, our cross-modal pre-training approach effectively integrates consistency cues from both modalities, achieving a top-1 accuracy of 97.8%.
@@
Table 3: Ablation studies on the major components of CCL-SLR.
@@ 4.4 Ablation Study
In this section, we conduct several ablation studies to verify the effectiveness of key components in our proposed framework on the MSASL [48] dataset.
@@
Cross-Modal Contrastive Learning Loss. As shown in Table 4, we first investigate the impact of the cross-modal contrastive learning loss $\mathcal{L}_{cross}$. After removing $\mathcal{L}_{cross}$ from CCL-SLR, the top-1 accuracy significantly drops by about 5%. The reason is that $\mathcal{L}_{cross}$ plays a crucial role in our framework by explicitly and effectively aligning the feature spaces between the two modalities, which implicitly helps the model extract cross-modal consistency cues. Additionally, Fig. 4 demonstrates the heatmaps generated by fine-tuning RGB encoder from pre-training with and without $\mathcal{L}_{cross}$. Especially, performing cross-modal contrastive learning during pretraining allows the RGB branch to learn semantic information from the pose modality. This significantly enhances the model’s attention to non-manual features (yellow and red areas) such as facial expressions and body movements.
Figure 4: Visualizations of heatmaps by Grad-CAM [35]. Top: raw frames; Middle: heatmaps of the fine-tuning results of the RGB encoder from pre-training without $\mathcal{L}_{cross}$; Bottom: heatmaps of the fine-tuning results of the RGB encoder from pre-training with $\mathcal{L}_{cross}$. All samples are from the NMFs-CSL [15] dataset.
Pre-Training Data Scale. In Fig. 5, we examine the impact of the pre-training data scale on the performance of our proposed framework. The experimental results reveal that the performance exhibits a steady improvement as the proportion of the pre-training data scale increases, demonstrating that our model can effectively leverage large-scale unlabeled multi-modal sign language data to enhance its representation capability, highlighting the scalability of our approach to large-scale sign language data.
Figure 5: Hyperparameter selection on CCL-SLR performance. We systematically investigate the impact of pre-training data scale, mask probability $\alpha$ in MPM, and number of pseudo-positive labels $k$ in SPM.
Motion-Preserving Masking. The influence of MPM is explored by sequentially changing the mask probability parameter $\alpha$ within $[0.0,1.0]$. As shown in Fig.5, the performance reaches its peak at $\alpha=0.2$, indicating that moderate masking effectively mitigates static bias while preserving essential dynamic information. However, an excessively high mask ratio can lead to performance degradation. Visualization results of MPM are provided in the Appendix.
Semantic Positive Mining. We investigate the sensitivity of hyperparameter $k$ to our method, which denotes the number of mined semantic positives per modality. As illustrated in Fig. 5, experimental results demonstrate that either extremely small or large values of $k$ lead to performance degradation. An overly large $k$ will introduce noise during contrastive learning, negatively impacting the representation capability of the model. Based on these empirical findings, we set $k=10$ as the default value.
@@
Table 6: Ablation studies on different pre-training modalities.
Figure 6: t-SNE visualizations of feature embeddings. We sample 30 sign words from the SLR500 dataset and visualize the feature embeddings extracted by CCL-SLR and its baseline, respectively.
Memory Banks. Unlike the vanilla MoCo v2 [3] framework, our pipeline maintains two memory banks ($\mathcal{M}_{g}$ and $\mathcal{M}_{f}$) in our pipeline: one for pseudo-labeling and the other for contrastive learning. Intuitively, maintaining a unique memory bank is structurally more concise. However, according to previous work [9], features before projection are more generally applicable, while features after projection are specialized for the instance discrimination task in contrastive learning. As shown in the Table 5, our experimental results further confirm that directly using $\mathcal{M}_{f}$ for pseudo-labeling leads to a slight performance degradation.
@@ 5 Conclusion
In this paper, we investigate the challenge of sign language recognition through cross-modal consistency learning. We propose CCL-SLR, a novel self-supervised multi-modal pre-training framework that enhances sign representation by leveraging consistency information across RGB and pose modalities. Additionally, our framework incorporates two key innovations: MPM, which enables motion-aware regions extraction from RGB videos to tackle the problem of redundant features, and SPM, which effectively addresses the semantic collision problem in contrastive learning. We conduct comprehensive evaluations on four widely used ISLR datasets, where CCL-SLR consistently achieves impressive performance, thereby validating the effectiveness of our proposed approach.